OpenStudy (anonymous):

the problem was m = 5, passing through P(0, 7) and according to my book the answer was wrong. can someone show how to write an equation for this

OpenStudy (anonymous):

y - 7 = 5(x -0)

OpenStudy (anonymous):

y - 7 = 5( x - 0)

OpenStudy (anonymous):

what form is your question requiring

OpenStudy (anonymous):

or slope intercept?

OpenStudy (anonymous):

which is y = 5x + 7 or standard form:5x - y = -7 in case you need it